# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v3

@@ 2 Challenges
$-\boldsymbol{\Theta}^{-1}+\boldsymbol{S}+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\|\boldsymbol{\Theta}\|_{1},$ (2.1)
@@
$-\boldsymbol{\Theta}_{D}^{-1}+({1}/{2})(\boldsymbol{\Theta}\boldsymbol{S}+\boldsymbol{S}\boldsymbol{\Theta})+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\|\boldsymbol{\Theta}_{X}\|_{1}.$ (2.3)
@@ 3.1 ACCORD loss and estimator
For the associated risk $R(\boldsymbol{\Omega}):=\operatorname{\bf E}L(\boldsymbol{\Omega},X)=-\log\det\boldsymbol{\Omega}_{D}+\frac{1}{2}\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{\Sigma}^{*})$ with the essential domain $\operatorname{\bf dom}R=\{(m_{ij})\in\mathbb{R}^{p\times p}:m_{ii}>0,~{}i=1,\dotsc,p\}\cap\mathbb{R}^{p\times p}$ on which $R$ is finite, the following theorem holds (details are in the Supplemental Material).
@@
$-\boldsymbol{\Omega}_{D}^{-1}+\boldsymbol{\Omega}\boldsymbol{S}+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\left\|\boldsymbol{\Omega}\right\|_{1},$ (3.3)
@@
For the ACCORD estimator (3.2) to be well-defined, a solution to (3.3) needs to be unique. A matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ $(p\geq n)$ is said to have columns in general position if the affine span of any $n$ points $\{s_{1}X_{i_{1}},\cdots,s_{n}X_{i_{n}}\}$, for arbitrary signs $s_{1},\cdots,s_{n}\in\{-1,1\}$, does not contain any element of {±$\{\pm X_{i}:i\neq i_{1},\cdots,i_{n}\}$. General position occurs almost surely if $\boldsymbol{X}$ is drawn from a continuous distribution [38].
@@ 3.2 ACCORD-FBS algorithm
$\begin{split}\omega^{(t+1)}_{ii}&=(y-\tau_{t}\lambda+\sqrt{(y-\tau_{t}\lambda)^{2}+4\tau_{t}})/2,\quad y=\omega^{(t)}_{ii}-\tau_{t}[\boldsymbol{\Omega}^{(t)}\boldsymbol{S}]_{ii}\\ \omega^{(t+1)}_{ij}&=S_{\tau_{t}\lambda}(\omega^{(t)}_{ij}-\tau_{t}[\boldsymbol{\Omega}^{(t)}\boldsymbol{S}]_{ij}),\quad i\neq j,\end{split}$ (3.5)
@@
Input: sample covariance $\boldsymbol{S}\in\mathbb{R}^{p\times p}$, minimum step size $1/L$, initial step size $\tau_{0}$, line search parameter $10<\beta<$ 1, initial $\boldsymbol{\Omega}^{(0)}$
@@
         $\nabla g(\boldsymbol{\Omega}^{(t)})\leftarrow\boldsymbol{\Omega}^{(t)}\boldsymbol{S}$triangleright Use Algorithm 2 for HP-ACCORD
@@
Assume the condition for Theorem 3.2, and that the iterate sequence $\{\boldsymbol{\Omega}^{(t)}\}$ is generated by Algorithm 1 with the step size sequence $\{\tau_{t}\}$ satisfying either i) $\beta=1$, $\tau_{0}\in[\underline{\tau},\bar{\tau}]$, where $0<\underline{\tau}\leq\bar{\tau}<2/L$, or ii) $10<\beta<$ 1. Then, the objective value sequence $\{f(\boldsymbol{\Omega}^{(t)})\}$ converges to the minimum $f^{\star}=f(\hat{\boldsymbol{\Omega}})$ monotonically. Furthermore, if the initial iterate $\boldsymbol{\Omega}^{(0)}$ is chosen such that $f(\boldsymbol{\Omega}^{(0)})\leq f(\boldsymbol{I}_{p})=\operatorname{\bf tr}(\boldsymbol{S})/2+\lambda p$, then the following holds.
$\begin{split}f(\boldsymbol{\Omega}^{(t)})-f^{\star}&\leq\left(\frac{1}{1+2a\sigma}\right)^{t}[f(\boldsymbol{\Omega}^{(0)})-f^{\star}],\quad t\geq 0,\\ \|\boldsymbol{\Omega}^{(t)}-\hat{\boldsymbol{\Omega}}\|_{F}&\leq\left(\frac{1}{\sqrt{1+2a\sigma}}\right)^{t-1}\left(1+\frac{1}{a\sigma\sqrt{1+{1}/{(2a\sigma)}}}\right)\sqrt{{[f(\boldsymbol{\Omega}^{(0)})-f^{\star}]}/{a}},\quad t\geq 1,\end{split}$ (3.6)
@@
Splitting (3.4) is crucial for establishing linear convergence in that the global Lipschitz constant $L<\infty$ of $\nabla g$ exists, and $01/L>$ 0 serves as the lower bound of the step size that guarantees the descent property. Following CONCORD-ISTA [29], we may also split $f$ into $\tilde{g}(\boldsymbol{\Omega})=-\log\det\boldsymbol{\Omega}_{D}+({1}/{2})\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})$ and $\tilde{h}(\boldsymbol{\Omega})=\lambda\|\boldsymbol{\Omega}\|_{1}$ and call it ACCORD-ISTA. The problem with this splitting is that $\nabla\tilde{g}$ is not globally Lipschitz. The consequence is that the resulting step sizes satisfying the descent condition can be arbitrarily small, contributing to slow convergence of the iterates; see §2 and §5.
@@
In Algorithm 1, the computational complexity of each iteration is $O(np^{2})$, with the bottleneck being the gradient computation $\nabla g(\boldsymbol{\Omega}^{(t)})\leftarrow(1/n)(\boldsymbol{\Omega}^{(t)}\boldsymbol{X}^{T})\boldsymbol{X}$; see Appendix C for comparison of the complexity with other methods. In fact, the required number of arithmetic operations can be further reduced by exploiting the sparsity of the iterate $\boldsymbol{\Omega}^{(t)}$. Moreover, matrix multiplication of the type $\boldsymbol{\Omega}^{(t)}\boldsymbol{X}^{T}$ can be distributed among multiple computational nodes, allowing the algorithm to be more efficient in HPC environments.
@@ 3.3 HP-ACCORD: HPC implementation of ACCORD
         send $\boldsymbol{\Omega}_{k}$ to node $k-j$ and recv $\boldsymbol{\Omega}_{k+j}$ from node $k+j$
@@ 4.2 Edge selection and sign consistency
[…]
@@ 5.1 Linear convergence of ACCORD-FBS
We provide empirical evidences of the guaranteed descent property and linear convergence exhibited by ACCORD-FBS (Algorithm 1). The merit of the novel operator splitting utilized by the latter is demonstrated by comparing it with ACCORD-ISTA (see the last paragraph of §3.2). Firstly, we generated two Erdos-Renyi graphs of size $p\!=\!1000$ with a sparsity level of $15\%$. To construct a ground truth precision matrix, we employed the following procedure. The edge weights were selected from a uniform distribution on $[0.5,1]$ and their signs were flipped with a probability of $0.5$. To ensure symmetry and positive definiteness, the resulting matrix was added to its transpose and its diagonal entries are set to be 1.5 times the absolute sum of the off-diagonal entries of the corresponding rows. This matrix was scaled by pre- and post-multiplying a diagonal matrix such that all the diagonal entries are equal to one. Finally, variation among the diagonal entries was introduced by performing another pre- and post-multiplications with a diagonal matrix with uniformly distributed entries on $[1,\sqrt{3}]$. Using this precision matrix, multivariate Gaussian data with a sample size of $n=500$ were generated.
@@ 5.2 Scalability of HP-ACCORD
We next investigate the scalability of HP-ACCORD (§3.3) using simulated data of dimension $p$ up to one million. At this scale, even simulating multivariate Gaussian samples becomes a nontrivial task [41]. We took a similar approach to the numerical experiments in Li et al. [23]: generate a sparse $p\times p$ unit lower triangular matrix $\boldsymbol{L}$ taking values in $[-1,1]$ and compute $\boldsymbol{y}=\boldsymbol{L}^{-T}\boldsymbol{x}$, $\boldsymbol{x}\sim N(\mathbf{0},\boldsymbol{I}_{p})$, by backsubstitution so that $\boldsymbol{y}$ follows $N(0,(\boldsymbol{L}\boldsymbol{L}^{T})^{-1})$. The location of off-diagonal non-zero entries of $\boldsymbol{L}$ were uniformly selected so that the graph implied by the precision matrix $\boldsymbol{L}\boldsymbol{L}^{T}$ had an average degree of $10.3$ for each $p$, and the maximum degree ranged from 42 to 66.
@@
We report the results in Figure 2. Since the memory of a single KNL node (96GB) was not enough for the computation with dimensions greater than or equal to $100,000$, an appropriate number of nodes that can handle the input size data was used for HP-ACCORD, showcasing the scalability of the algorithm. The supercomputing center imposed a 48-hour restriction on the running time of a single job. Thus we only report accurate timing for the processes finished within this limit. In panel (a), the regularization parameter $\lambda$ was set large enough so that the resulting precision matrix estimate $\hat{\boldsymbol{\Theta}}$ becomes diagonal. In panel (b), $\lambda$ was adjusted so that the numbers of nonzero entries of $\hat{\boldsymbol{\Theta}}$ have a similar scale to their true precision matrices. Note that fastclime had to use the time near the limit to estimate a precision matrix to compute estimate for two $\lambda$’s for $p=10,000$, and it could not run with data that has $p=30,000$ and above due to the memory limit. Also, for the both panel, we could observe that HP-ACCORD outperforms BigQUIC in terms of computation time, even when only a single node is employed. By adopting multiple computational nodes, HP-ACCORD was able to deliver the estimate within the time budget in all scenarios. On the contrary, BigQUIC could not finish the process on time when the dimension exceeds 100,000 and the $\lambda$ is chosen so that the estimate is non-diagonal. Internally, BigQUIC tries to find a permuted partition $\{B_{1},\cdots,B_{k}\}$ of the $p$ coordinates using a graph clustering algorithm [18, 9] so that most of the coordinate updates are performed in the diagonal blocks $(B_{1},B_{1}),\cdots,(B_{k},B_{k})$. However, the size of each partition is limited due to the memory limit of a single computational node (about 20,000 in the Nurion environment in which each node has 96GB of memory), and clustering with such a limited block size bares many off-block diagonal edges in our simulated data. In consequence, the conjugate gradient method employed to compute the off-diagonal blocks are hardly skipped, causing the algorithm to run extremely slowly. Even when $\lambda$ is so large that the estimate should be diagonal, BigQUIC failed to complete the computation if the dimension reached one million.
@@ 5.3 Edge detection and estimation error
Table 1 reports the area under the precision-recall curve (AUPRC) and number of edges selected. In terms of AUPRC (and Matthews correlation coefficient; see Supplementary Material B), ACCORD performed slightly better than or similarly to other methods. The trend of the precision-recall curves did not particularly vary among replications. In all cases, the penalty selected from each selection method yielded reasonable number of selected edges. We could observe that BIC-type methods with the extended term drastically reduces false positive (FP) edges in the expense of few true positive (TP) edges. Table 2 reports the total squared error of the estimated precision matrix $\boldsymbol{\Theta}$ and its reparameterization $\boldsymbol{\Omega}$ by ACCORD and CONCORD, along with their debiased refit. Compared to CONCORD, ACCORD estimates clearly showed better results for both $\hat{\boldsymbol{\Theta}}$ and $\hat{\boldsymbol{\Omega}}$ in terms of the estimation error. The debiased refit also improves the estimation performance. For more details about the impact of the debiased refit, see Supplementary Material B.
@@ 5.4 Estimation error dependency on precision matrix
We further provide numerical results that illustrate how the structures and values of the partial correlation matrix can affect the estimation error of ACCORD. Motivated by experiments in Ravikumar et al. [33], we conducted simulation studies under two types of $\boldsymbol{\Theta}^{*}$: chain and star graphs. In both types, we let $\theta_{ii}=1,$ implying that $\boldsymbol{\Omega}^{*}=\boldsymbol{\Theta}^{*}$. In the chain graph setting, we fix $p=120$ and let $\theta_{ij}=\rho>0$ for $|i-j|=1$ and $\theta_{ij}=0$ for $|i-j|\geq 2.$ Thus, the maximum number of non-zero entries per row $d$ is fixed to $2$. With $\rho$ increasing, the $\kappa_{\Omega^{*}}$ and $\kappa_{\Gamma^{*}}$ also increase and the maximum irrepresentability constant $\alpha$ in 4.1 decreases. Note that for $\rho\geq 0.5$, 4.1 holds for any $\alpha\in[0,1)$. In the star graph setting, $d-1$ nodes among $p=200$ nodes are connected only to a single hub node, and there are no other connections. For all the connected edges, partial correlation entries are set as $2.5/(d-1)$ so that $\kappa_{\Omega^{*}}$ remains invariant with $d$. For sufficiently large $d$, both $\kappa_{\Gamma^{*}}$ and $\alpha$ remain nearly constant. In both settings, an ACCORD estimate is computed with $n$ Gaussian samples drawn from the given $\boldsymbol{\Theta}^{*}$, and the penalty coefficient $\lambda$ is set to be proportional to $\sqrt{1/n}$ as suggested by the theory.
In Figure 3 we provide the maximum estimation error observed with various sample sizes $n$ for the two types of graphs. In the chain graph setting (panel (a)), the rates of the estimation error were similar for $\rho\leq 0.4$, while its magnitude increased with $\rho$. For $\rho=0.5$, the estimator could not recover the chain graph structure, and increasing the sample size did not further improve the estimation error, indicating the need for the irrepresentability condition. In the star graph setting (panel (b)), changing $d$ did not significantly affect the estimation error. This observation is consistent with Theorem 4.2.

### v4

@@ 2 Challenges
$\boldsymbol{\Theta}^{-1}+\boldsymbol{S}+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\|\boldsymbol{\Theta}\|_{1},$ (2.1)
@@
$\boldsymbol{\Theta}_{D}^{-1}+({1}/{2})(\boldsymbol{\Theta}\boldsymbol{S}+\boldsymbol{S}\boldsymbol{\Theta})+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\|\boldsymbol{\Theta}_{X}\|_{1}.$ (2.3)
@@ 3.1 ACCORD loss and estimator
For the associated risk $R(\boldsymbol{\Omega}):=\operatorname{\bf E}L(\boldsymbol{\Omega},X)=-\log\det\boldsymbol{\Omega}_{D}+\frac{1}{2}\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{\Sigma}^{*})$ with the essential domain $\operatorname{\bf dom}R=\{(m_{ij})\in\mathbb{R}^{p\times p}:m_{ii}>0,~i=1,\dotsc,p\}\cap\mathbb{R}^{p\times p}$ on which $R$ is finite, the following theorem holds (details are provided in the Supplemental Material).
@@
$\boldsymbol{\Omega}_{D}^{-1}+\boldsymbol{\Omega}\boldsymbol{S}+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\left\|\boldsymbol{\Omega}\right\|_{1},$ (3.3)
@@
For the ACCORD estimator (3.2) to be well-defined, a solution to (3.3) needs to be unique. A matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ $(p\geq n)$ is said to have columns in general position if the affine span of any $n$ points $\{s_{1}X_{i_{1}},\cdots,s_{n}X_{i_{n}}\}$, for arbitrary signs $s_{1},\cdots,s_{n}\in\{-1,1\}$, does not contain any element of $\{\pm X_{i}:i\neq i_{1},\cdots,i_{n}\}$. General position occurs almost surely if $\boldsymbol{X}$ is drawn from a continuous distribution [38].
@@ 3.2 ACCORD-FBS algorithm
$\begin{split}\omega^{(t+1)}_{ii}&=(y-\tau_{t}\lambda+\sqrt{(y-\tau_{t}\lambda)^{2}+4\tau_{t}})/2,\qquad y=\omega^{(t)}_{ii}-\tau_{t}[\boldsymbol{\Omega}^{(t)}\boldsymbol{S}]_{ii}\\ \omega^{(t+1)}_{ij}&=S_{\tau_{t}\lambda}(\omega^{(t)}_{ij}-\tau_{t}[\boldsymbol{\Omega}^{(t)}\boldsymbol{S}]_{ij}),\qquad i\neq j,\end{split}$ (3.5)
@@
Input: sample covariance $\boldsymbol{S}\in\mathbb{R}^{p\times p}$, minimum step size $1/L$, initial step size $\tau_{0}$, line search parameter $0<\beta<1$, initial $\boldsymbol{\Omega}^{(0)}$
@@
    $\nabla g(\boldsymbol{\Omega}^{(t)})\leftarrow\boldsymbol{\Omega}^{(t)}\boldsymbol{S}$ $\triangleright$ Use Algorithm 2 for HP-ACCORD
@@
[…]
$\begin{split}f(\boldsymbol{\Omega}^{(t)})-f^{\star}&\leq\left(\frac{1}{1+2a\sigma}\right)^{t}[f(\boldsymbol{\Omega}^{(0)})-f^{\star}],\qquad t\geq 0,\\ \|\boldsymbol{\Omega}^{(t)}-\hat{\boldsymbol{\Omega}}\|_{F}&\leq\left(\frac{1}{\sqrt{1+2a\sigma}}\right)^{t-1}\left(1+\frac{1}{a\sigma\sqrt{1+{1}/{(2a\sigma)}}}\right)\sqrt{{[f(\boldsymbol{\Omega}^{(0)})-f^{\star}]}/{a}},\qquad t\geq 1,\end{split}$ (3.6)
@@
Splitting (3.4) is crucial for establishing linear convergence in that the global Lipschitz constant $L<\infty$ of $\nabla g$ exists, and $1/L>0$ serves as the lower bound of the step size that guarantees the descent property. Following CONCORD-ISTA [29], we may also split $f$ into $\tilde{g}(\boldsymbol{\Omega})=-\log\det\boldsymbol{\Omega}_{D}+({1}/{2})\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})$ and $\tilde{h}(\boldsymbol{\Omega})=\lambda\|\boldsymbol{\Omega}\|_{1}$ and call it ACCORD-ISTA. The problem with this splitting is that $\nabla\tilde{g}$ is not globally Lipschitz. The consequence is that the resulting step sizes satisfying the descent condition can be arbitrarily small, contributing to slow convergence of the iterates; see §2 and §5.
@@
In Algorithm 1, the computational complexity of each iteration is $O(np^{2})$, with the bottleneck being the gradient computation $\nabla g(\boldsymbol{\Omega}^{(t)})\leftarrow(1/n)(\boldsymbol{\Omega}^{(t)}\boldsymbol{X}^{T})\boldsymbol{X}$; see Supplementary Material C for comparison of the complexity with other methods. In fact, the required number of arithmetic operations can be further reduced by exploiting the sparsity of the iterate $\boldsymbol{\Omega}^{(t)}$. Moreover, matrix multiplication of the type $\boldsymbol{\Omega}^{(t)}\boldsymbol{X}^{T}$ can be distributed among multiple computational nodes, allowing the algorithm to be more efficient in HPC environments.
@@ 3.3 HP-ACCORD: HPC implementation of ACCORD
    send $\boldsymbol{\Omega}_{k}$ to node kjk-j and recv $\boldsymbol{\Omega}_{k+j}$ from node $k+j$
@@ 4.2 Edge selection and sign consistency
where $\sigma$ and $\delta$ are as defined in Theorem 4.2, then the perfect sign recovery event $\{\operatorname{sign}(\omega_{ij}^{*})=\operatorname{sign}(\hat{\omega}_{ij})$ for all (i,j)}(i,j)\} occurs with a probability no smaller than $1-p^{-(\tau-2)}$.
@@ 5.1 Linear convergence of ACCORD-FBS
We provide empirical evidences of the guaranteed descent property and linear convergence exhibited by ACCORD-FBS (Algorithm 1). The merit of the novel operator splitting utilized by the latter is demonstrated by comparing it with ACCORD-ISTA (see the last paragraph of §3.2). Firstly, we generated two Erdos-Renyi graphs of size p=1000p\!=\!1000 with a sparsity level of $15\%$. To construct a ground truth precision matrix, we employed the following procedure. The edge weights were selected from a uniform distribution on $[0.5,1]$ and their signs were flipped with a probability of $0.5$. To ensure symmetry and positive definiteness, the resulting matrix was added to its transpose and its diagonal entries are set to be 1.5 times the absolute sum of the off-diagonal entries of the corresponding rows. This matrix was scaled by pre- and post-multiplying a diagonal matrix such that all the diagonal entries are equal to one. Finally, variation among the diagonal entries was introduced by performing another pre- and post-multiplications with a diagonal matrix with uniformly distributed entries on $1,\sqrt{3}]$. Using this precision matrix, multivariate Gaussian data with a sample size of $n=500$ were generated.
@@ 5.2 Scalability of HP-ACCORD
We next investigate the scalability of HP-ACCORD (§3.3) using simulated data of dimension $p$ up to one million. At this scale, even simulating multivariate Gaussian samples becomes a nontrivial task [41]. We took a similar approach to the numerical experiments in Li et al. [23]: generate a sparse $p\times p$ unit lower triangular matrix $\boldsymbol{L}$ taking values in [1,1][-1,1] and compute $\boldsymbol{y}=\boldsymbol{L}^{-T}\boldsymbol{x}$, $\boldsymbol{x}\sim N(\mathbf{0},\boldsymbol{I}_{p})$, by backsubstitution so that $\boldsymbol{y}$ follows $N(0,(\boldsymbol{L}\boldsymbol{L}^{T})^{-1})$. The location of off-diagonal non-zero entries of $\boldsymbol{L}$ were uniformly selected so that the graph implied by the precision matrix $\boldsymbol{L}\boldsymbol{L}^{T}$ had an average degree of $10.3$ for each $p$, and the maximum degree ranged from 42 to 66.
@@
We report the results in Figure 2. Since the memory of a single KNL node (96GB) was not enough for the computation with data dimensions greater than or equal to $100,000$, an appropriate number of nodes that can handle the input size data was used for HP-ACCORD, showcasing the scalability of the algorithm. The supercomputing center imposed a 48-hour restriction on the running time of a single job. Thus we only report accurate timing for the processes finished within this limit. In panel (a), the regularization parameter $\lambda$ was set large enough so that the resulting precision matrix estimate $\hat{\boldsymbol{\Theta}}$ becomes diagonal. In panel (b), $\lambda$ was adjusted so that the numbers of nonzero entries of $\hat{\boldsymbol{\Theta}}$ have a similar scale to their true precision matrices. Note that fastclime had to use the time near the limit to estimate a precision matrix to compute estimate for two $\lambda$’s for $p=10,000$, and it could not run on data with $p=30,000$ or higher due to time and memory limitations. Also, for the both panel, we could observe that HP-ACCORD outperforms BigQUIC in terms of computation time, even when only a single node is employed. By adopting multiple computational nodes, HP-ACCORD was able to deliver the estimate within the time budget in all scenarios. On the contrary, BigQUIC failed to finish the process on time when the dimension exceeded 100,000 and the $\lambda$ was chosen to yield a non-diagonal estimator. Internally, BigQUIC tries to find a permuted partition $\{B_{1},\cdots,B_{k}\}$ of the $p$ coordinates using a graph clustering algorithm [18, 9] so that most of the coordinate updates are performed in the diagonal blocks $(B_{1},B_{1}),\cdots,(B_{k},B_{k})$. However, the size of each partition is limited due to the memory limit of a single computational node (about 20,000 in the Nurion environment in which each node has 96GB of memory), and clustering with such a limited block size bares many off-block diagonal edges in our simulated data. In consequence, the conjugate gradient method employed to compute the off-diagonal blocks are hardly skipped, causing the algorithm to run extremely slowly. Even when $\lambda$ is so large that the estimate should be diagonal, BigQUIC failed to complete the computation if the dimension reached one million.
@@ 5.3 Edge detection and estimation error
Table 1 reports the area under the precision-recall curve (AUPRC) and number of edges selected. In terms of AUPRC (and Matthews correlation coefficient; see Supplementary Material B), ACCORD performed slightly better than or similarly to other methods. The trend of the precision-recall curves did not particularly vary among replications. In all cases, the penalty selected from each selection method yielded reasonable number of selected edges. We could observe that BIC-type methods with the extended term drastically reduces false positive (FP) edges in the expense of few true positive (TP) edges.
Table 2 reports the total squared error of the estimated precision matrix $\boldsymbol{\Theta}$ and its reparameterization $\boldsymbol{\Omega}$ by ACCORD and CONCORD, along with their debiased refit. Compared to CONCORD, ACCORD estimates clearly showed better results for both $\hat{\boldsymbol{\Theta}}$ and $\hat{\boldsymbol{\Omega}}$ in terms of the estimation error. The debiased refit also improves the estimation performance. For more details about the impact of the debiased refit, see Supplementary Material B.
@@ 5.4 Estimation error dependency on precision matrix
We further provide numerical results that illustrate how the structures and values of the partial correlation matrix can affect the estimation error of ACCORD. Motivated by experiments in Ravikumar et al. [33], we conducted simulation studies under two types of $\boldsymbol{\Theta}^{*}$: chain and star graphs. In both types, we let $\theta_{ii}=1,$ implying that $\boldsymbol{\Omega}^{*}=\boldsymbol{\Theta}^{*}$. In the chain graph setting, we fix $p=120$ and let $\theta_{ij}=\rho>0$ for |ij|=1|i-j|=1 and $\theta_{ij}=0$ for $|i-j|\geq 2.$ Thus, the maximum number of non-zero entries per row $d$ is fixed to $2$. With $\rho$ increasing, the $\kappa_{\Omega^{*}}$ and $\kappa_{\Gamma^{*}}$ also increase and the maximum irrepresentability constant $\alpha$ in 4.1 decreases. Note that for $\rho\geq 0.5$, 4.1 holds for any $\alpha\in[0,1)$. In the star graph setting, d1d-1 nodes among $p=200$ nodes are connected only to a single hub node, and there are no other connections. For all the connected edges, partial correlation entries are set as 2.5/(d1)2.5/(d-1) so that $\kappa_{\Omega^{*}}$ remains invariant with $d$. For sufficiently large $d$, both $\kappa_{\Gamma^{*}}$ and $\alpha$ remain nearly constant. In both settings, an ACCORD estimate is computed with $n$ Gaussian samples drawn from the given $\boldsymbol{\Theta}^{*}$, and the penalty coefficient $\lambda$ is set to be proportional to $\sqrt{1/n}$ as suggested by the theory.
In Figure 3 we provide the maximum estimation error observed with various sample sizes $n$ for the two types of graphs. In the chain graph setting (panel (a)), the rates of the estimation error were similar for $\rho\leq 0.4$, while its magnitude increased with $\rho$. For $\rho=0.5$, the estimator could not recover the chain graph structure, and increasing the sample size did not further improve the estimation error, indicating the necessity of the irrepresentability condition. In the star graph setting (panel (b)), changing $d$ did not significantly affect the estimation error. This observation is consistent with Theorem 4.2.